# ChatBug: A Common Vulnerability of Aligned LLMs Induced by Chat Templates

### v1

@@ Abstract
In this paper, we investigate how chat templates affect safety alignment of LLMs. We identify a common vulnerability, named ChatBug, that is introduced by chat templates. Our key insight to identify ChatBug is that the chat templates provide a rigid format that need to be followed by LLMs, but not by users. Hence, a malicious user may not necessarily follow the chat template when prompting LLMs. Instead, malicious users could leverage their knowledge of the chat template and accordingly craft their prompts to bypass safety alignments of LLMs. We develop two attacks to exploit the ChatBug vulnerability. We demonstrate that a malicious user can exploit the ChatBug vulnerability of eight state-of-the-art (SOTA) LLMs and effectively elicit unintended responses from these models. Moreover, we show that ChatBug can be exploited by existing jailbreak attacks to enhance their attack success rates. We investigate potential countermeasures to ChatBug. Our results show that while adversarial training effectively mitigates the ChatBug vulnerability, the victim model incurs significant performance degradation. These results highlight the trade-off between safety alignment and helpfulness. Developing new methods for instruction tuning to balance this trade-off is an open and critical direction for future research [^1] Our code is available at https://github.com/uw-nsl/ChatBug..
ChatBug: A Common Vulnerability of Aligned LLMs
Induced by Chat Templates
[…]
WARNING: This paper contains model outputs that may be considered offensive.
[…]
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 (Achiam et al., 2023) and Llama-3 (Meta, 2024) have been increasingly used to empower conversational agents. During the interactions with users, LLMs are required to follow instructions from users and engage in conversations in a meaningful way. Standard techniques to enhance instruction-following capabilities include instruction tuning (Bakker et al., 2022; Wei et al., 2022) and preference tuning (Bai et al., 2022; Christiano et al., 2017; Ouyang et al., 2022a).
A common practice for instruction tuning and preference tuning is to structure data using chat templates OpenAI (2023). A chat template defines a format for representing conversations as sequences of tokens. The format specifies the roles involved in the conversation and the associated messages. Chat templates are shown to be effective in optimizing the LLMs’ performance (Jiang et al., 2023; OpenAI, 2023), allowing LLMs to generate coherent responses in their designated roles.
In addition to instruction-following capabilities, LLMs are also required to generate responses that are aligned with human values. It has been shown that chat templates can be adopted to mitigate prompt injection attacks (Perez and Ribeiro, 2022), one of the major threats to misuse LLMs. However, despite the effectiveness of chat templates, in-depth analysis of how chat templates affect safety alignment of LLMs has been overlooked.
@@
We show that any malicious user who has knowledge of the chat template can exploit the ChatBug vulnerability. This assumption is not restrictive, especially for open-source models whose chat templates are often publicly available. We develop two attacks, format mismatch attack and message overflow attack as illustrated in Figure 1, to exploit the vulnerability and elicit unsafe responses from LLMs. Specifically, the format mismatch attack modifies the default chat format, while the message overflow attack injects a sequence of tokens into the model’s reserved field. We present proof-of-concept for each attack to demonstrate its feasibility and effectiveness. We demonstrate the severity and pervasivity of the ChatBug vulnerability on eight LLMs (Vicuna (Chiang et al., 2023), Mistral (Jiang et al., 2023), Llama-2 (Touvron et al., 2023), Llama-3 (Meta, 2024), GPT-3.5 (OpenAI, 2023), Gemini (Google, 2023), Claude-2.1 (Anthropic, 2023), and Claude-3 (Anthropic, 2024)). We further show that existing jailbreak attacks such as GCG Zou et al. (2023), GPTFuzzer (Yu et al., 2023), and ArtPrompt (Jiang et al., 2024) can exploit the ChatBug vulnerability to increase their attack success rates.
@@ 2 Related Work
Prior work focuses on how the design of prompts impacts model performance, such as the prompting language (Gonen et al., 2023), order of few-shot examples (Lu et al., 2022), example choices (Xiang et al., 2024), and prompt formatting (Sclar et al., 2024). This paper, which investigates how chat templates affects safety alignment of LLMs, is parallel to the aforementioned studies.
@@
Extensive efforts have been made to align LLMs with human values. Standard techniques include supervised fine-tuning (Bakker et al., 2022; Wei et al., 2022), preference tuning (Bai et al., 2022; Christiano et al., 2017; Ouyang et al., 2022a), and red-teaming (Dinan et al., 2019; Ge et al., 2023). Despite these efforts, jailbreak attacks (Wei et al., 2023) pose a significant threat to misuse of LLMs. Jailbreak attacks can be categorized into two classes based on how they bypass safety alignment. The first category designs attack prompts based on heuristics that rely on human experts (Deng et al., 2023; Huang et al., 2023; Li et al., 2023; Liu et al., 2023b; Jiang et al., 2024; Zeng et al., 2024). The second category utilizes optimization problems to search for prompts to jailbreak aligned LLMs. Gradient-based (Jones et al., 2023; Zou et al., 2023; Zhu et al., 2023), genetic algorithm-based (Liu et al., 2023a), and edit-based methods (Chao et al., 2023) have been developed to solve the optimization problems.
@@ 3.1 Preliminary Background
Let $\mathcal{M}$ represent an auto-regressive LLM whose vocabulary is denoted as $\mathcal{V}$. Given an input represented by a sequence of tokens of length $n$, denoted as $x_{1:n}=x_{1},\ldots,x_{n}$, the LLM predicts a probability distribution $p_{\mathcal{M}}(\cdot|x_{1:n})$ over the vocabulary $\mathcal{V}$. Then the LLM samples the next token $x_{n+1}\in\mathcal{V}$ according to a predefined decoding strategy (e.g., greedy or beam search (Wu et al., 2016)) and probability distribution $p_{\mathcal{M}}$. Appending token $x_{n+1}$ to the sequence $x_{1:n}$ and iteratively applying the procedure for next token generation as described above yield text generation by the LLM. This process continues until a stopping criterion is met, such as reaching the maximum generation length or generating an end-of-sequence (EOS) token.
@@
Instruction tuning is the critical step to enable a pre-trained LLM to follow instructions from users. Such processes include supervised fine-tuning, and/or reinforcement-learning from human feedback (RLHF) Ouyang et al. (2022b). Different from pre-training, instruction tuning employs a chat template to structure data in the form of (multi-turn) conversations.
An example of chat template, named ChatML (OpenAI, 2023), is presented in Table 1. The template defines a format for representing conversations using a sequence of tokens. It starts by segmenting a conversation involving multiple turns into individual turns, where the segments are separated by a set of special control tokens, denoted as beginning-of-turn (BOT) and end-of-turn (EOT) tokens. For each turn, the template organizes the dialogue by role control tokens (e.g., ‘user’ and ‘assistant’) and their respective messages. These pieces of tokens within a single turn are delineated by BOT and EOT tokens. Note that the BOT and EOT tokens are different from beginning-of-sequence (BOS) and end-of-sequence (EOS) tokens, which are used for both base LLMs and aligned LLMs to delimit the beginning and end of a sequence.
@@ 3.2 Chat Templates Induce a Common Vulnerability: ChatBug
Although chat templates are effective in fine-tuning LLMs to function as conversational agents, we highlight that they introduce a common vulnerability, named ChatBug, to LLMs. A malicious user, with knowledge of the format predefined by the chat template, could exploit this vulnerability to elicit harmful responses from victim LLMs by crafting its queries. Our key insight is that the chat templates pre-define rigid formats that should be followed by LLMs, but not by users. For example, the malicious user could craft its query by appending the role of LLM and the beginning of desired harmful response at the end, as illustrated in Appendix B. Consequently, the malicious user tricks victim LLMs to complete the harmful responses provided by the user, instead of continuing the conversation in its intended role. We note that it is not a restrictive assumption to suggest that malicious users could have access to the knowledge of chat templates, as these are often publicly available (HuggingFace, 2024). As we will discuss later, exploiting this vulnerability does not rely on white-box (Qi et al., 2024) or grey-box access (Zou et al., 2023) to the victim model.
@@
Table 1: Format defined by the ChatML chat template (OpenAI, 2023): <|im_start|> and <|im_end|> are BOT and EOT tokens. user and assistant are role control tokens. The corresponding messages are ‘How are you’ and ‘I am doing well!’.
@@ 4.1 Format Mismatch Attack
For instance, a malicious user may omit all control tokens including BOT, EOT, and role control tokens when prompting the victim LLMs. The insight behind the format mismatch attack is that the format specified by the chat template is not mandatory for users to follow. Since many LLMs may not verify whether user queries match the format required by the chat template, these modifications may induce a different interpretation of input queries to victim LLMs, leading to harmful or unintended responses. An example of the format mismatch attack can be found in Appendix B.
Figure 2: We denote the sequence of tokens crafted using the format mismatch attack as $\hat{x}_{1:n}$ and the sequence of tokens designed according to the chat template as $x_{1:n}$. This figure presents how the ratio $\frac{P_{\mathcal{M}}(\cdot|\hat{x}_{1:n})}{P_{\mathcal{M}}(\cdot|x_{1:n})}$ evolves at each step $n$, with the results averaged over 50 instructions. The format mismatch attack significantly increases the probability of generating the desired harmful response.
@@
Figure 3: This figure presents the probability of generating the desired harmful response when the number of overflow tokens varies from 0 to 9, with the results averaged over 50 instructions. Note that the user does not launch message overflow attack when the number of overflow tokens is zero. The results show that the probability of generating the desired harmful response increases as the user overflows more tokens.
@@ 4.2 Message Overflow Attack
where $s$ is the overflow message from the malicious user. Consequently, the victim LLMs are tricked to complete the harmful response based on their auto-regressive generation capabilities, instead of continuing the conversation with users in their designated roles. An example of the message overflow attack is presented in Appendix B.
@@ 4.3 ChatBug Boosts Jailbreak Attacks
Existing ailbreak attacks Jiang et al. (2024); Yu et al. (2023); Zou et al. (2023) elicit unintended responses from victim LLMs using deliberately designed prompts, which correspond to user message $m$ in the chat templates. By exploiting the ChatBug vulnerability and launching jailbreak attacks, malicious users could jointly tamper with the chat template via the two attacks mentioned above and messages within a conversation with victim LLMs. Consequently, the malicious users can effectively boost the probability of generating unintended responses from victim LLMs.
@@ 5.1 Experimental Setup
We evaluate ChatBug on eight LLMs including both open-source and closed-source models. For open-source models, we select Vicuna (7B-v1.5) (Chiang et al., 2023), Llama-2 (7B-Chat) (Touvron et al., 2023), Llama-3 (8B-Instruct) (Meta, 2024), Mistral (7B-Instruct v0.2) (Jiang et al., 2023). For closed-source models, we consider GPT-3.5 developed by OpenAI [^2] We use Microsoft AZure OpenAI service for the experiment in our work: https://azure.microsoft.com/en-us/products/ai-services/openai-service, Claude-2.1 and Claude-3 (Opus) from Anthropic (Anthropic, 2023, 2024), as well as Gemini (Pro) from Google (Google, 2023).
@@
We assess the severity of ChatBug using a metric named attack success rate (ASR), which is defined as follows:
@@
Following Zou et al. (2023) and Yu et al. (2023), we calculate ASR using two approaches:
  • Refusal Response Matching (ASR-R). This approach collects a set of refusal responses (e.g., “Sorry, I cannot …") and verifies whether the response generated by LLM matches any of them. An LLM generated response is considered harmful if it does not align with any of the refusal responses.
  • Moderator Assessment (ASR-M). We utilize a pretrained LLM, Llama-Guard-2333https://huggingface.co/meta-llama/Meta-Llama-Guard-2-8B fine-tuned from Llama 3, as a moderator to evaluate whether a response is harmful.
@@
Table 2: This table summarizes the ASR-R and ASR-M of Direct Instruct (baseline) and attacks that exploit the ChatBug vulnerability in open-source LLMs. The results show that an attacker can effectively bypass safety alignment of LLMs by exploiting the ChatBug vulnerability. We have excluded results of Mismatch-V on Vicuna model since they use the same chat template.
@@
Table 3: This table summarizes the ASR-R and ASR-M of Direct Instruct (baseline) and attacks that exploit the ChatBug vulnerability in closed-source LLMs. The results show that an attacker can effectively bypass safety alignment of LLMs by exploiting the ChatBug vulnerability. We have excluded results of format mismatch attacks on Gemini and Claude models since their APIs do not support features to execute such attacks.
@@
  • Mismatch-C. An attacker launches the format mismatch attack by replacing the default chat template with ChatML developed by OpenAI (2023).
  • Mismatch-V. In this case, an attacker initiates the format mismatch attack by replacing the default chat template with the one used by Vicuna (Chiang et al., 2023).
  • Overflow-S. An attacker launches the message overflow attack. For any query, the overflow inserts a fixed prefix ‘Sure, here is’ for the desired response from the victim model.
@@
  • Overflow-FS. This is an extension of Overflow-L, where an attacker launches the message overflow attack and utilizes an uncensored LLM to generate the prefixes. In our experiments, we choose Wizard-Vicuna-13B-Uncensored [^4] https://huggingface.co/cognitivecomputations/Wizard-Vicuna-13B-Uncensored as the uncensored LLM, and use in-context-learning prompting technique to generate multiple prefixes that carry affirmative semantics automatically. We generate 5 prefixes by default. The prompt to generate attack prefixes is in Figure 5 in Appendix.

### v2

@@ Abstract
In this paper, we investigate how chat templates affect safety alignment of LLMs. We identify a common vulnerability, named ChatBug, that is introduced by chat templates. Our key insight to identify ChatBug is that the chat templates provide a rigid format that need to be followed by LLMs, but not by users. Hence, a malicious user may not necessarily follow the chat template when prompting LLMs. Instead, malicious users could leverage their knowledge of the chat template and accordingly craft their prompts to bypass safety alignments of LLMs. We study two attacks to exploit the ChatBug vulnerability. Additionally, we demonstrate that the success of multiple existing attacks can be attributed to the ChatBug vulnerability. We show that a malicious user can exploit the ChatBug vulnerability of eight state-of-the-art (SOTA) LLMs and effectively elicit unintended responses from these models. Moreover, we show that ChatBug can be exploited by existing jailbreak attacks to enhance their attack success rates. We investigate potential countermeasures to ChatBug. Our results show that while adversarial training effectively mitigates the ChatBug vulnerability, the victim model incurs significant performance degradation. These results highlight the trade-off between safety alignment and helpfulness. Developing new methods for instruction tuning to balance this trade-off is an open and critical direction for future research.
Codehttps://github.com/uw-nsl/ChatBug
Extended versionhttps://arxiv.org/abs/2406.12935
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 (Achiam et al. 2023) and Llama-3 (Meta 2024) have been increasingly used to empower conversational agents. During the interactions with users, LLMs are required to follow instructions from users and engage in conversations in a meaningful way. Standard techniques to enhance instruction-following capabilities include instruction tuning (Bakker et al. 2022; Wei et al. 2022) and preference tuning (Bai et al. 2022; Christiano et al. 2017; Ouyang et al. 2022a).
A common practice for instruction tuning and preference tuning is to structure data using chat templates (OpenAI 2023). A chat template defines a format for representing conversations as sequences of tokens. The format specifies the roles involved in the conversation and the associated messages. Chat templates are shown to be effective in optimizing the LLMs’ performance (Jiang et al. 2023; OpenAI 2023), allowing LLMs to generate coherent responses in their designated roles.
In addition to instruction-following capabilities, LLMs are also required to generate responses that are aligned with human values. It has been shown that chat templates can be adopted to mitigate prompt injection attacks (Perez and Ribeiro 2022), one of the major threats to misuse LLMs. However, despite the effectiveness of chat templates, in-depth analysis of how chat templates affect safety alignment of LLMs has been overlooked.
@@
We show that any malicious user who has knowledge of the chat template can exploit the ChatBug vulnerability. This assumption is not restrictive, especially for open-source models whose chat templates are often publicly available. We study two attacks, format mismatch attack and message overflow attack as illustrated in Figure 1, to exploit the vulnerability and elicit unsafe responses from LLMs. Specifically, the format mismatch attack modifies the default chat format, while the message overflow attack injects a sequence of tokens into the model’s reserved field. Specifically, the message overflow attack unifies multiple existing attacks (Zou et al. 2023; Vega et al. 2023; Andriushchenko, Croce, and Flammarion 2024), whose success can all be attributed to the ChatBug vulnerability. We demonstrate the severity and pervasivity of the ChatBug vulnerability on eight LLMs (Vicuna (Chiang et al. 2023), Mistral (Jiang et al. 2023), Llama-2 (Touvron et al. 2023), Llama-3 (Meta 2024), GPT-3.5 (OpenAI 2023), Gemini (Google 2023), Claude-2.1 (Anthropic 2023), and Claude-3 (Anthropic 2024)). We further show that existing jailbreak attacks such as GCG (Zou et al. 2023), GPTFuzzer (Yu, Lin, and Xing 2023), and ArtPrompt (Jiang et al. 2024b) can exploit the ChatBug vulnerability to increase their attack success rates.
@@ 2 Related Work
Prior work focuses on how the design of prompts impacts model performance, such as the prompting language (Gonen et al. 2023), order of few-shot examples (Lu et al. 2022), example choices (Xiang et al. 2024), and prompt formatting (Sclar et al. 2024). This paper, which investigates how chat templates affects safety alignment of LLMs, is parallel to the aforementioned studies.
@@
Extensive efforts have been made to align LLMs with human values. Standard techniques include supervised fine-tuning (Bakker et al. 2022; Wei et al. 2022), preference tuning (Bai et al. 2022; Christiano et al. 2017; Ouyang et al. 2022a), and red-teaming (Dinan et al. 2019; Ge et al. 2023). Despite these efforts, jailbreak attacks (Wei, Haghtalab, and Steinhardt 2023) pose a significant threat to misuse of LLMs. Jailbreak attacks can be categorized into two classes based on how they bypass safety alignment. The first category designs attack prompts based on heuristics that rely on human experts (Deng et al. 2023; Huang et al. 2023; Li et al. 2023b; Liu et al. 2023b; Jiang et al. 2024b; Zeng et al. 2024; Lin et al. 2024). The second category utilizes optimization problems to search for prompts to jailbreak aligned LLMs. Gradient-based (Jones et al. 2023; Zou et al. 2023; Zhu et al. 2023), genetic algorithm-based (Liu et al. 2023a), and edit-based methods (Chao et al. 2023) have been developed to solve the optimization problems.
@@ 3.1 Preliminary Background
Let $\mathcal{M}$ represent an auto-regressive LLM whose vocabulary is denoted as $\mathcal{V}$. Given an input represented by a sequence of tokens of length $n$, denoted as $x_{1:n}=x_{1},\ldots,x_{n}$, the LLM predicts a probability distribution $p_{\mathcal{M}}(\cdot|x_{1:n})$ over the vocabulary $\mathcal{V}$. Then the LLM samples the next token $x_{n+1}\in\mathcal{V}$ according to a predefined decoding strategy (e.g., greedy or beam search (Wu et al. 2016)) and probability distribution $p_{\mathcal{M}}$. Appending token $x_{n+1}$ to the sequence $x_{1:n}$ and iteratively applying the procedure for next token generation as described above yield text generation by the LLM. This process continues until a stopping criterion is met, such as reaching the maximum generation length or generating an end-of-sequence (EOS) token.
@@
Instruction tuning is the critical step to enable a pre-trained LLM to follow instructions from users. Such processes include supervised fine-tuning, and/or reinforcement-learning from human feedback (RLHF) (Ouyang et al. 2022b). Different from pre-training, instruction tuning employs a chat template to structure data in the form of (multi-turn) conversations. An example of chat template, named ChatML (OpenAI 2023), is presented in Table 1. The template defines a format for representing conversations using a sequence of tokens. It starts by segmenting a conversation involving multiple turns into individual turns, where the segments are separated by a set of special control tokens, denoted as beginning-of-turn (BOT) and end-of-turn (EOT) tokens. For each turn, the template organizes the dialogue by role control tokens (e.g., ‘user’ and ‘assistant’) and their respective messages. These tokens within a single turn are delineated by BOT and EOT tokens. Note that the BOT and EOT tokens are different from beginning-of-sequence (BOS) and end-of-sequence (EOS) tokens, which are used for both base LLMs and aligned LLMs to delimit the beginning and end of a sequence.
@@ 3.2 Chat Templates Induce a Common Vulnerability: ChatBug
Although chat templates are effective in fine-tuning LLMs to function as conversational agents, we highlight that they introduce a common vulnerability, named ChatBug, to LLMs. A malicious user, with knowledge of the format predefined by the chat template, could exploit this vulnerability to elicit harmful responses from victim LLMs by crafting its queries. Our key insight is that the chat templates pre-define rigid formats that should be followed by LLMs, but not by users. For example, the malicious user could craft its query by appending the role of LLM and the beginning of desired harmful response at the end, as illustrated in Appendix B of our extended version (Jiang et al. 2024a). Consequently, the malicious user tricks victim LLMs to complete the harmful responses provided by the user, instead of continuing the conversation in its intended role. We note that it is not a restrictive assumption to suggest that malicious users could have access to the knowledge of chat templates, as these are often publicly available (HuggingFace 2024). As we will discuss later, exploiting this vulnerability does not rely on white-box (Qi et al. 2024) or grey-box access (Zou et al. 2023) to the victim model.
@@
Table 1: Format defined by the ChatML chat template (OpenAI 2023): <|im_start|> and <|im_end|> are BOT and EOT tokens. user and assistant are role control tokens. The corresponding messages are ‘How are you’ and ‘I am doing well!’.
@@ 4.1 Format Mismatch Attack
For instance, a malicious user may omit all control tokens including BOT, EOT, and role control tokens when prompting the victim LLMs. The insight behind the format mismatch attack is that the format specified by the chat template is not mandatory for users to follow. Since many LLMs may not verify whether user queries match the format required by the chat template, these modifications may induce a different interpretation of input queries to victim LLMs, leading to harmful or unintended responses. An example of the format mismatch attack can be found in Appendix B of our extended version (Jiang et al. 2024a). We note that the format mismatch attack alters the chat template, distinguishing it from the setup of jailbreak attacks, which focus on manipulating the prompt message but following the standard chat template (Liu et al. 2023a; Li et al. 2023b).
Figure 2: This figure shows how the ratio $\frac{P_{\mathcal{M}}(\cdot|\hat{x}_{1:n})}{P_{\mathcal{M}}(\cdot|x_{1:n})}$ evolves at each decoding step $n$ (i.e., the number of response tokens), with the results averaged over 50 instructions. The format mismatch attack significantly increases the probability of generating the desired harmful response.
@@
Figure 3: This figure presents the probability of generating the desired harmful response when the number of overflow tokens varies from 0 to 9, averaged over 50 instructions. Note that the user does not launch message overflow attack when the number of overflow tokens is zero. The results show that the probability of generating the desired harmful response increases as the user overflows more tokens.
@@ 4.2 Message Overflow Attack
where $s$ is the overflow message from the malicious user. Consequently, the victim LLMs are tricked to complete the harmful response based on their auto-regressive generation capabilities, instead of continuing the conversation with users in their designated roles. An example of the message overflow attack is presented in Appendix B of our extended version (Jiang et al. 2024a). The message overflow attack provides a unified framework for existing attacks (Zou et al. 2023; Vega et al. 2023; Andriushchenko, Croce, and Flammarion 2024), which adopt seemingly distinct attack strategies. We highlight that the success of these attacks can all be attributed to the ChatBug vulnerability.
@@ 4.3 ChatBug Boosts Jailbreak Attacks
Existing ailbreak attacks (Jiang et al. 2024b; Yu, Lin, and Xing 2023; Zou et al. 2023) elicit unintended responses from victim LLMs using deliberately designed prompts, which correspond to user message $m$ in the chat templates. By exploiting the ChatBug vulnerability and launching jailbreak attacks, malicious users could jointly tamper with the chat template via the two attacks mentioned above and messages within a conversation with victim LLMs. Consequently, the malicious users can effectively boost the probability of generating unintended responses from victim LLMs.
@@ 5.1 Experimental Setup
We evaluate ChatBug on eight LLMs including both open-source and closed-source models. For open-source models, we select Vicuna (7B-v1.5) (Chiang et al. 2023), Llama-2 (7B-Chat) (Touvron et al. 2023), Llama-3 (8B-Instruct) (Meta 2024), Mistral (7B-Instruct v0.2) (Jiang et al. 2023). For closed-source models, we consider GPT-3.5 developed by OpenAI [^1] We use Microsoft AZure OpenAI service for the experiment in our work: https://azure.microsoft.com/en-us/products/ai-services/openai-service, Claude-2.1 and Claude-3 (Opus) from Anthropic (Anthropic 2023, 2024), as well as Gemini (Pro) from Google (Google 2023).
@@
We assess the severity of ChatBug using a metric named attack success rate (ASR), which is defined as
@@
Following Zou et al. (2023) and Yu, Lin, and Xing (2023), we calculate ASR using two approaches:
  • Refusal Response Matching (ASR-R). This approach collects a set of refusal responses (e.g., “Sorry, I cannot …”) and verifies whether the response generated by LLM matches any of them. An LLM generated response is considered harmful if it does not align with any of the refusal responses.
  • Moderator Assessment (ASR-M). We utilize a pretrained LLM, Llama-Guard-$2$https://huggingface.co/meta-llama/Meta-Llama-Guard-2-8B fine-tuned from Llama 3, as a moderator to evaluate whether a response is harmful.
@@
Table 2: This table summarizes the ASR-R and ASR-M of Direct Instruct (baseline) and attacks that exploit the ChatBug in open-source LLMs. The results show that an attacker can effectively bypass safety alignment of LLMs by exploiting the ChatBug vulnerability. We have excluded results of Mismatch-V on Vicuna model since they use the same chat template.
@@
Table 3: This table summarizes the ASR-R and ASR-M of Direct Instruct (baseline) and attacks that exploit the ChatBug vulnerability in closed-source LLMs. The results show that an attacker can effectively bypass safety alignment of LLMs by exploiting the ChatBug vulnerability. We have excluded results of format mismatch attacks on Gemini and Claude models since their APIs do not support features to execute such attacks. We detail how to launch format mismatch attacks against commercial LLMs with API access in Appendix A.2 of our extended version (Jiang et al. 2024a).
@@
  • Mismatch-C. An attacker launches the format mismatch attack by replacing the default chat template with a complex template ChatML developed by OpenAI (2023).
  • Mismatch-V. An attacker initiates the format mismatch attack by replacing the default chat template with a simple template used by Vicuna (Chiang et al. 2023).
  • Overflow-S. An attacker launches the message overflow attack. For any query, the overflow inserts a fixed prefix ‘Sure, here is’ for the desired response.
@@
  • Overflow-FS. This is an extension of Overflow-L, where an attacker launches the message overflow attack and utilizes an uncensored LLM to generate the prefixes. In our experiments, we choose Wizard-Vicuna-13B-Uncensored [^3] https://huggingface.co/cognitivecomputations/Wizard-Vicuna-13B-Uncensored as the uncensored LLM, and use in-context-learning prompting technique to generate multiple prefixes that carry affirmative semantics automatically. We generate 5 prefixes by default. The prompt to generate attack prefixes is in Figure 7 of our extended version (Jiang et al. 2024a).